# Improving LLM Agents with Reinforcement Learning on Cryptographic CTF Challenges

### v1

@@ 4.2 Reinforcement Learning
w/ Hints $0.38$ $0.00$ +0.13 +0.00
w/o Hints $0.38$ $0.00$ +0.13 +0.00
Curriculum $0.38$ $0.00$ +0.13 +0.00
5 Disscussion
Executing model-generated code introduces a spectrum of safety and security concerns, particularly in reinforcement learning settings where agents are incentivized to solve tasks efficiently. In our experiments, an MCP-based Python REPL server [21] was used to enable persistent tool-use across multiple calls in a single trajectory. This design preserved the variable context between invocations, simulating a lightweight computation environment with memory.
However, this setup revealed concrete risks. We observed that when the agent generated code with excessive memory requirements—such as attempting to enumerate all $6$-character ASCII strings into a list—the REPL server unconditionally executed it. As a result, the underlying container crashed due to memory exhaustion. This highlights a critical vulnerability in any naive integration of tool-augmented agents with unbounded execution environments.
@@ 6 Conclusion
In this study, we demonstrated that reinforcement learning, specifically Guided Reinforcement Prompt Optimization (GRPO), significantly improves the cryptographic problem-solving abilities of LLM agents. By introducing and employing the random-crypto benchmark, we systematically evaluated how curriculum-based training and hint incorporation influence agent performance. Our results confirm that structured reinforcement learning substantially enhances not only the accuracy and efficiency of solutions but also the robustness and correctness of tool usage. Furthermore, we provided evidence that improvements from training generalize beyond the original dataset, highlighting genuine skill acquisition rather than mere dataset-specific optimization.
@@ Acknowledgements
The research was supported by the Hungarian National Research, Development and Innovation Office within the framework of the Thematic Excellence Program 2021 – National Research Sub programme: “Artificial intelligence, large networks, data security: mathematical foundation and applications” and the Artificial Intelligence National Laboratory Program (MILAB). We would also like to thank GitHub and neptune.ai for providing us academic access. Special thanks to Alex Hornyai for helping us validate some of the harder crypto challenges.
Appendix 0.A Additional Dataset Details
@@
Web Crypto JWT ‘none‘ algorithm, Weak cookie encryption, Broken key exchange, Insecure session token
@@
Appendix 0.B Prompts used in the paper
@@
The following prompt was used to instruct the language model to generate a short narrative for each cryptographic challenge. The variable CIPHER is replaced with the name of the cipher used in the instance, and VULNERABILITY is an optional string describing the nature of the weakness being exploited.
Prompt 1: Challenge Story Generation Prompt Write a very short ($\leq$ 30 words) back-story for a crypto challenge that uses the CIPHER cipher VULNERABILITY. Keep it very brief without unnecesary information. Return only the text between these XML tags: <challenge>…</challenge> Insert the placeholder <CIPHER> verbatim where the ciphertext will go. Do not write anything outside the tags.
@@
This prompt instructs the language model to perform tool-augmented reasoning using the execute_python tool. The placeholder MCP_TOOL_LIST is replaced at runtime with a list of available tools, and QUESTION is the current question to solve.

### v2

@@ 4.3 Generalization
w/ Hints 0.18 0.08 +0.11 +0.03
w/o Hints $0.08$ $0.05$ +0.01 +0.00
Curriculum 0.18 $0.07$ +0.11 +0.02
We further evaluated the trained agents on the AICrypto MCQ benchmark [22], a 135-question multiple-choice dataset focused exclusively on cryptography. Each question offered 4–8 possible answers and spanned a wide range of topics, from purely theoretical items to computational tasks requiring scripting to be solved. The model had never been trained on multiple-choice data; however, no filtering or prompt adaptation was applied; we used the same tool-augmented prompt format as in Random-Crypto training. As shown in Table 5, the curriculum-trained agent achieved the largest improvement over the baseline, increasing Pass@8 by +0.15 and Maj@8 by +0.07, despite the format shift. The base model and the other training variants tended to hallucinate the wrong answer after reasoning. This suggests that curriculum fine-tuning may improve an agent’s adaptability to unseen task formats, balancing both reasoning and computational tool use across heterogeneous cryptographic problems.
Table 5: RL training’s effect on performance on the external AICrypto MCQ benchmark.
After Training Improvement
Training Type Pass@8 Maj@8 Pass@8 Maj@8
w/ Hints $0.04$ $0.00$ +0.00 +0.00
w/o Hints $0.03$ $0.00$ 0.01-0.01 +0.00
Curriculum $0.19$ $0.07$ +0.15 +0.07
5 Discussion
Executing code generated by LLMs introduces a spectrum of safety and security concerns, particularly in Reinforcement Learning settings where agents are incentivized to solve tasks efficiently. In our experiments, an MCP-based Python REPL server [27] was used to enable persistent tool-use across multiple calls in a single trajectory. This design preserved variable state across multiple invocations, simulating a lightweight computation environment with memory.
However, this setup revealed concrete risks. We observed that when the agent generated code with excessive memory requirements—such as attempting to enumerate all $6$-character ASCII strings into a list, the REPL server unconditionally executed it. As a result, the underlying container crashed due to memory exhaustion. This highlights a critical vulnerability in any naive integration of tool-augmented agents with unbounded execution environments.
@@ 6 Conclusion
This study demonstrates that Reinforcement Learning can provide Large Language Model (LLM) agents with durable and transferable problem-solving abilities in security-critical domains. By leveraging the Random-Crypto benchmark, we fine-tuned a tool-augmented agent using Group Relative Policy Optimization (GRPO) and observed significant improvements in cryptographic reasoning, procedural tool use, and cross-domain generalization.
Our results indicate that RL is not merely optimizing outputs; it enables agents to acquire robust internal routines for reasoning and tool interaction. The procedural structure of Random-Crypto played a key role: by offering abundant, verifiable, and tool-centric challenges, it supports scalable, curriculum-aligned agent training. The gains observed on the AICrypto MCQ benchmark, despite the model never encountering multiple-choice data during training, reinforce that RL-fine-tuning can instill transferable problem-solving routines. In particular, the curriculum-trained agent’s advantage here may indicate improved flexibility in adapting tool-augmented reasoning strategies to novel task formats.
Looking ahead, this framework opens the door to agent-centric Reinforcement Learning in cybersecurity and beyond—where competence, adaptability, and safe tool use become foundational design objectives.
@@ Acknowledgements
The research was supported by the Hungarian National Research, Development and Innovation Office within the framework of the Thematic Excellence Program 2021 – National Research Sub programme: “Artificial intelligence, large networks, data security: mathematical foundation and applications” and the Artificial Intelligence National Laboratory Program (MILAB). We would also like to thank GitHub and neptune.ai for providing us with academic access. Special thanks to Alex Hornyai for helping us validate some of the harder crypto challenges.
Appendix A Additional Dataset Details
@@
Web Crypto JWT ‘none’ algorithm, Weak cookie encryption, Broken key exchange, Insecure session token
@@
Appendix B Prompts used in the paper
@@
The following prompt was used to instruct the LLM to generate a short narrative for each cryptographic challenge. The variable CIPHER is replaced with the name of the cipher used in the instance, and VULNERABILITY is an optional string describing the nature of the weakness being exploited.
Prompt 1: Challenge Story Generation Prompt Write a very short ($\leq$ 30 words) back-story for a crypto challenge that uses the CIPHER cipher VULNERABILITY. Keep it very brief without unnecessary information. Return only the text between these XML tags: <challenge>…</challenge> Insert the placeholder <CIPHER> verbatim where the ciphertext will go. Do not write anything outside the tags.
@@
This prompt instructs the LLM to perform tool-augmented reasoning using the execute_python tool. The placeholder MCP_TOOL_LIST is replaced at runtime with a list of available tools, and QUESTION is the current question to solve.